# Unifying Multimodal Large Language Model Capabilities and Modalities via Model Merging

### v1

@@ 7 Conclusion
Acknowledgments
We would like to thank Chi Chen and Shixin Jiang for their assistance in collecting datasets related to modality merging. We are also grateful to the Qwen Team, OpenGVLab, and the developers of the open-source models used in this work for their generous contributions. Additionally, we thank the creators of VLMEvalKit and LMMs-Eval for providing essential evaluation tools.
@@ Appendix A Theoretical Proofs
Assumption A.2 has been formulated in various ways throughout the literature. For convex conditions, [28] proposed the bound $4\mathcal{C}\bigl{(}\mathcal{L}(\Theta)-\mathcal{L}(\Theta_{*})\bigr{)}+2\sigma^{2}$, where $\Theta_{*}$ is the global minimizer and $\sigma^{2}\triangleq\mathbb{E}\left[\left\|g(\Theta_{*})\right\|^{2}\right]$. Similarly, for non-convex settings, [39] introduced the expected smoothness assumption as $2A\bigl{(}\mathcal{L}(\Theta)-\mathcal{L}^{\inf}\bigr{)}+B\,\|\nabla\mathcal{L}(\Theta)\|^{2}+C$, for some $A,B,C\geq 0$ and $\forall\Theta\in\mathbb{R}^{d}$.
@@
Denote the merged task vector as $\boldsymbol{\tau}_{m}=\sum_{j}\alpha_{j}\boldsymbol{\tau}_{j}$, and the task vector of task $i$ as $\boldsymbol{\tau}_{i}$. Let $\mathcal{L}_{i}$ be the loss for task $i$. The difference in loss is bounded by the following: $\left\|\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{m})-\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{i})\right\|\leq\textstyle\sum_{j}\Psi_{j}^{i}\cdot\left\|\boldsymbol{\tau}_{j}\right\|,$ (7) where $\Psi_{j}^{i}$ is the coefficient capturing how sensitive task $i$’s loss is to parameter updates from task $j$ defined as: $\Psi_{j}^{i}=\begin{cases}\mathcal{C}_{i}\cdot|\alpha_{j}|,&\text{if }j\neq i\\ \mathcal{C}_{i}\cdot|1-\alpha_{j}|,&\text{if }j=i.\end{cases}$ (8)
@@
$\left\|\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{m})-\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{i})\right\|$ $\leq\mathcal{C}_{i}\cdot\left\|\boldsymbol{\tau}_{m}-\boldsymbol{\tau}_{i}\right\|$ (9)
$=\mathcal{C}_{i}\cdot\textstyle\left\|\sum_{j}\alpha_{j}\,\boldsymbol{\tau}_{j}-\boldsymbol{\tau}_{i}\right\|$ (10)
$\leq\mathcal{C}_{i}\cdot(\textstyle\sum_{j\neq i}\alpha_{j}\cdot\left\|\boldsymbol{\tau}_{j}\right\|+|1-\alpha_{i}|\cdot\|\boldsymbol{\tau}_{i}\|).$ (11)
@@
Applying Eq. 12 to Eq. 11, we obtain:
$\left\|\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{m})-\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{i})\right\|\leq\textstyle\sum_{j}\Psi_{j}^{i}\cdot\left\|\boldsymbol{\tau}_{j}\right\|.$ (13)
@@
Let $\boldsymbol{\theta}_{0}\in\mathbb{R}^{d}$ be the initial parameter. For each task $i$, after $T$ gradient steps with constant learning rate $\eta$, we denote the task vector as $\boldsymbol{\tau}_{i}=\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{0}$. Considering that most methods can be viewed as extensions of linear combinations of task vectors, let $\boldsymbol{\tau}_{m}=\sum_{j}\alpha_{j}\boldsymbol{\tau}_{j}$ denote the merged vector. The loss on task $i$ is denoted by $\mathcal{L}_{i}(\Theta)$, which is $\mathcal{C}_{i}$-Lipschitz continuous. Assuming a constraint on the second moment of the gradient, then: $\left\|\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{m})-\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{i})\right\|\leq\mathcal{O}(\eta T)$ (14) This indicates that both the learning rate and iterations influence model merging results.
@@
$\boldsymbol{\tau}_{j}=\textstyle\sum_{t=0}^{T-1}-\eta\cdot g_{t}(\Theta),$ (15)
@@
$\mathbb{E}[\|\boldsymbol{\tau}_{j}\|^{2}]$ $=\mathbb{E}[\|\textstyle\sum_{t=0}^{T-1}-\eta\cdot g_{t}(\Theta)\|^{2}]$ (16)
@@
$\mathbb{E}\left[\|\boldsymbol{\tau}_{j}\|\right]$ $\leq\sqrt{\mathbb{E}\left[\|\boldsymbol{\tau}_{j}\|^{2}\right]}$ (20)
@@
Applying Eq. 20 to Lemma A.3, we obtain:
$\left\|\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{m})-\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{i})\right\|\leq\underbrace{(\textstyle\sum_{j}\Psi_{j}^{i})\cdot\mathcal{G}}_{\text{coefficient}}\cdot\eta T.$ (22)
@@
$\left\|\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{m})-\mathcal{L}_{i}(\boldsymbol{\theta}_{0}+\boldsymbol{\tau}_{i})\right\|\leq\mathcal{O}(\eta T).$ (23)
Appendix B Fine-Tuning Influence on Model Merging
To demonstrate the sensitivity of model merging to task vectors $\boldsymbol{\tau}_{i}$ (i.e., parameter changes between fine-tuned models and the base model), we conduct experiments using the standard CLIP-ViT model merging [35], following the fine-tuning setup of FusionBench [73]. The Adam optimizer is employed with a fixed learning rate of 1e-5, running for 4000 training steps with a batch size of 32. Models are saved every 500 iterations and evaluated for accuracy on the test dataset, as illustrated in Fig. 5. Convergence is generally observed around 3000 iterations.
Additionally, we evaluate four classical merging methods at various fine-tuning stages, reporting their average accuracy in Fig. 6. The results indicate that increasing the number of fine-tuning steps does not consistently enhance merging performance. Instead, performance typically improves initially before declining. This finding motivates our derivation of Theorem 3.1, which demonstrates that both the learning rate and the number of iterations affect model merging performance. MLLM training is typically organized by full passes over the data (epochs), rather than discrete iteration steps. Accordingly, we set the number of epochs to 1 and reduce the learning rate to limit parameter changes. This keeps the fine-tuned models close to the base model in parameter space while still yielding improvements on specific tasks.
@@ Appendix B Fine-Tuning Influence on Model Merging
Appendix C Challenges of MLLMs Merging Benchmark
@@
Training and evaluation of MLLMs are developed independently rather than being split from the same dataset. (i) Recent benchmarks [50, 42, 22] often assess models’ comprehensive abilities through pre-defined choice questions, with each benchmark emphasizing different nuanced aspects. (ii) Domain-specific training data is also proprietary and confidential. Consequently, models demonstrate varied capabilities based on their training foundations. For example, LLaVA [49] excels in conversational visual reasoning, while InstructBLIP [18] performs better on traditional short-answer VQA tasks. These differences present challenges for developing a unified benchmark suitable for multi-task model merging.
@@
Traditional public datasets provide rich visual-specific capabilities but lack instruction-following formats. On the other hand, instruction-following data generated by models like GPT-4 [2] often lack task-specific characteristics. This creates tension between these two essential aspects of model performance.
@@
Most publicly available models are already instruction-tuned versions that have undergone extensive training on diverse datasets, including open-source repositories, curated purchased datasets, and private collections. Further SFT often yields diminishing returns, as models may already be overfit to the common datasets used for training [33].
Appendix D Details of Modality Setting
Following DAMC [7], we select modality encoders and connectors paired with Vicuna-7B-v1.5 to train separate models using bi-modal data across three modalities (i.e., vision, audio and video). Additional details are presented in Table 8. The training approach consists of two phases: an alignment stage where only connector parameters are trainable, and a fine-tuning stage where we tune all connector and language model parameters. During fine-tuning, we apply LoRA with a rank of 128 across all linear modules within the LLM. For our merging strategy, we preserve each modality’s unique encoder and connector components while merging only the language model parameters, enabling the model to process inputs from all three modalities simultaneously.
@@ Appendix D Details of Modality Setting
Vision CLIP-ViT-L-336px [66] MLP LCS 558K [49] LLaVA-mixed 665K [48] LLaVA-1.5 [48]
@@
Audio BEATs-Iter3+ [9] Q-Former WaveCaps 400K [60] OpenAQA filtered 350K [27] X-InstructBLIP [64]
@@
Video LanguageBind [98] MLP LCS 558K [49], Valley 702K [53] Video-ChatGPT 100K [54], LLaVA-mixed subset 140K [48] Video-LLaVA [47]
Appendix E Understanding the Task Vector
@@
WUDI Merging [14] substitutes the transpose of the task vector $\boldsymbol{\tau}_{l}$ for the input vector $\boldsymbol{x}_{l}$. We reconsider the update process of the task vector, which can be formulated as follows:
$\boldsymbol{\tau}_{l}^{k}$ $=\textstyle\sum_{t=1}^{\mathrm{T}}-\eta\cdot\frac{\partial{\mathcal{L}(\boldsymbol{\theta}^{t-1})}}{\partial{\boldsymbol{\theta}^{t-1}_{l,k}}}$ (24)
$=\textstyle\sum_{t=1}^{\mathrm{T}}-\eta\sum^{N}_{n=1}\frac{\partial\mathcal{L}(\boldsymbol{\theta}^{t-1})}{\partial(\boldsymbol{\theta}^{t-1}_{l,k}\boldsymbol{x}_{n,l}^{t-1})}\cdot\frac{\partial(\boldsymbol{\theta}^{t-1}_{l,k}\boldsymbol{x}_{n,l}^{t-1})}{\partial\boldsymbol{\theta}^{t-1}_{l,k}}$ (25)
@@
$=\underbrace{\textstyle\sum_{t=1}^{\mathrm{T}}-\eta\sum^{N}_{n=1}\frac{\partial\mathcal{L}(\boldsymbol{\theta}^{t-1})}{\partial(\boldsymbol{\theta}^{t-1}_{l,k}\boldsymbol{x}_{n,l}^{t-1})}}_{\text{coefficient}}\cdot(\boldsymbol{x}_{n,l}^{t-1})^{\top},$ (26)
@@
where $\boldsymbol{\tau}_{l}^{k}$ denotes the task vector of neuron $k$ in linear layer $l$, and $\boldsymbol{\theta}^{t-1}_{l,k}$ represents the parameters of neuron $k$ in linear layer $l$ at time $t-1$. Each neuron in the linear layer can be interpreted as a weighted sum of input vectors across training iterations, with gradients serving as coefficients.
Appendix F Broader Impacts

### v2

@@ Appendix A Theoretical Proofs
Assumption A.2 has been formulated in various ways throughout the literature. For convex conditions, Gower et al. (2019) proposed the bound $4\mathcal{C}\bigl(\mathcal{L}(\Theta)-\mathcal{L}(\Theta_{*})\bigr)+2\sigma^{2}$, where $\Theta_{*}$ is the global minimizer and $\sigma^{2}\triangleq\mathbb{E}\left[\left\|g(\Theta_{*})\right\|^{2}\right]$. Similarly, for non-convex settings, Khaled & Richtárik (2023) introduced the expected smoothness assumption as $2A\bigl(\mathcal{L}(\Theta)-\mathcal{L}^{\inf}\bigr)+B\,\|\nabla\mathcal{L}(\Theta)\|^{2}+C$, for some $A,B,C\geq 0$ and $\forall\Theta\in\mathbb{R}^{d}$.
@@
Denote the merged task vector as $\bm{\tau}_{m}=\sum_{j}\alpha_{j}\bm{\tau}_{j}$, and the task vector of task $i$ as $\bm{\tau}_{i}$. Let $\mathcal{L}_{i}$ be the loss for task ii. The difference in loss is bounded by the following: $\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\textstyle\sum_{j}\Psi_{j}^{i}\cdot\left\|\bm{\tau}_{j}\right\|,$ (7) where $\Psi_{j}^{i}$ is the coefficient capturing how sensitive task $i$’s loss is to parameter updates from task $j$ defined as: $\Psi_{j}^{i}=\begin{cases}\mathcal{C}_{i}\cdot|\alpha_{j}|,&\text{if }j\neq i\\ \mathcal{C}_{i}\cdot|1-\alpha_{j}|,&\text{if }j=i.\end{cases}$ (8)
@@
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|$ $\leq\mathcal{C}_{i}\cdot\left\|\bm{\tau}_{m}-\bm{\tau}_{i}\right\|$ (9)
$=\mathcal{C}_{i}\cdot\textstyle\left\|\sum_{j}\alpha_{j}\,\bm{\tau}_{j}-\bm{\tau}_{i}\right\|$ (10)
$\leq\mathcal{C}_{i}\cdot(\sum_{j\neq i}\alpha_{j}\cdot\left\|\bm{\tau}_{j}\right\|$
$\quad+\,|1-\alpha_{i}|\cdot\|\bm{\tau}_{i}\|).$ (11)
@@
Applying Eq.˜12 to Eq.˜11, we obtain:
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\textstyle\sum_{j}\Psi_{j}^{i}\cdot\left\|\bm{\tau}_{j}\right\|.$ (13)
@@
Let $\bm{\theta}_{0}$ be the initial parameter. For each task $i$, after $T$ gradient steps with constant learning rate $\eta$, we denote the task vector as $\bm{\tau}_{i}=\bm{\theta}_{i}-\bm{\theta}_{0}$. Considering that most methods can be viewed as extensions of linear combinations of task vectors, let $\bm{\tau}_{m}=\sum_{j}\alpha_{j}\bm{\tau}_{j}$ denote the merged vector. The loss on task $i$ is denoted by $\mathcal{L}_{i}(\Theta)$, which is $\mathcal{C}_{i}$-Lipschitz continuous. Assuming a constraint on the second moment of the gradient, then: $\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\mathcal{O}(\eta T)$ (14) This indicates that both the learning rate and iterations influence model merging results.
@@
$\bm{\tau}_{j}=\textstyle\sum_{t=0}^{T-1}-\eta\cdot g_{t}(\Theta),$ (15)
@@
$\mathbb{E}[\|\bm{\tau}_{j}\|^{2}]$ $=\mathbb{E}[\|\textstyle\sum_{t=0}^{T-1}-\eta\cdot g_{t}(\Theta)\|^{2}]$ (16)
@@
$\mathbb{E}\left[\|\bm{\tau}_{j}\|\right]$ $\leq\sqrt{\mathbb{E}\left[\|\bm{\tau}_{j}\|^{2}\right]}$ (20)
@@
Applying Eq.˜20 to Lemma A.3, we obtain:
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\underbrace{(\textstyle\sum_{j}\Psi_{j}^{i})\cdot\mathcal{G}}_{\text{coefficient}}\cdot\eta T.$ (22)
@@
$\left\|\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{m})-\mathcal{L}_{i}(\bm{\theta}_{0}+\bm{\tau}_{i})\right\|\leq\mathcal{O}(\eta T).$ (23)
Appendix B Model Merging Benchmarks
B.1 Fine-Tuning Influence on Model Merging
To demonstrate the sensitivity of model merging to task vectors $\bm{\tau}_{i}$ (i.e., parameter changes between fine-tuned models and the base), we conduct experiments using the standard CLIP-ViT merging benchmark (Ilharco et al., 2023), following the fine-tuning setup of FusionBench (Tang et al., 2024a). We train with Adam (learning rate 1e-5) for 4,000 steps with a batch size of 32. Models are saved every 500 iterations and evaluated for accuracy on the test dataset, as illustrated in Fig.˜5. Across eight tasks, convergence typically occurs around 3,000 steps.
Additionally, we evaluate four classical merging methods at various fine-tuning stages, reporting their average accuracy in Fig.˜6. The results indicate that increasing the number of fine-tuning steps does not consistently enhance merging performance. Instead, performance typically improves initially before declining. This finding motivates our derivation of Theorem˜3.1, which demonstrates that both the learning rate and the number of iterations affect model merging performance. MLLM training is typically organized by full passes over the data (epochs), rather than discrete iteration steps. Accordingly, we set the number of epochs to 1 and reduce the learning rate to limit parameter changes. This keeps the fine-tuned models close to the base model in parameter space while still yielding improvements on specific tasks.
@@ B.1 Fine-Tuning Influence on Model Merging
B.2 Challenges of MLLMs Merging Benchmark
@@
Training and evaluation of MLLMs are developed independently rather than being split from the same dataset. (i) Recent benchmarks (Liu et al., 2024b; Li et al., 2024; Fu et al., 2024) often assess models’ comprehensive abilities through pre-defined choice questions, with each benchmark emphasizing different nuanced aspects. (ii) Domain-specific training data is also proprietary and confidential. Consequently, models demonstrate varied capabilities based on their training foundations. For example, LLaVA (Liu et al., 2023) excels in conversational visual reasoning, while InstructBLIP (Dai et al., 2023) performs better on traditional short-answer VQA tasks. These differences present challenges for developing a unified benchmark suitable for multi-task model merging.
@@
Public vision datasets provide strong task-specific supervision but rarely use instruction-following formats. Conversely, instruction data generated by models such as GPT-4 (Achiam et al., 2023) often lacks task-specific grounding. This mismatch creates a trade-off between instruction adherence and task expertise.
@@
Many publicly released models are already instruction-tuned on diverse sources, including open-source, licensed, and private datasets. Additional supervised fine-tuning (SFT) therefore yields diminishing returns and can further overfit models to widely used training distributions (Huang et al., 2025).
Appendix C Implementation Details
Checkpoint construction.
For InternVL2.5-1B-Instruct, we perform full fine-tuning with a learning rate of 4e-5 and a warmup ratio of 3e-2. For Qwen2-VL-7B-Base, we apply LoRA fine-tuning with a rank of 8, a learning rate of 1e-5, and a warmup ratio of 1e-1. Both models are trained for one epoch using a cosine learning rate scheduler. Different training strategies and scales help evaluate the generalizability of merging methods.
We follow prior work (Chen et al., 2024a) by pairing Vicuna-7B-v1.5 with modality-specific encoders and connectors. Separate models are trained using bi-modal data across three modalities: vision, audio, and video. Additional details are presented in Table˜8. The training approach consists of two phases: an alignment stage where only connector parameters are trainable, and a fine-tuning stage where we tune all connector and language model parameters. During fine-tuning, we apply LoRA with a rank of 128 across all linear modules within the LLM. For our merging strategy, we preserve each modality’s unique encoder and connector components while merging only the language model parameters, enabling the model to process inputs from all three modalities simultaneously.
@@ Appendix C Implementation Details
Vision CLIP-ViT-L-336px (Radford et al., 2021) MLP LCS 558K (Liu et al., 2023) LLaVA-mixed 665K (Liu et al., 2024a) LLaVA-1.5 (Liu et al., 2024a)
@@
Audio BEATs-Iter3+ (Chen et al., 2023) Q-Former WaveCaps 400K (Mei et al., 2024) OpenAQA filtered 350K (Gong et al., 2024) X-InstructBLIP (Panagopoulou et al., 2023)
@@
Video LanguageBind (Zhu et al., 2023) MLP LCS 558K (Liu et al., 2023), Valley 702K (Luo et al., 2023) Video-ChatGPT 100K (Maaz et al., 2024), LLaVA-mixed subset 140K (Liu et al., 2024a) Video-LLaVA (Lin et al., 2024)
Training data.
Following (Chen et al., 2024c), we collect a broader range of domain-specific data, divided into VQA, Geometry, Chart, OCR, and Grounding tasks. Specifically, for grounding tasks, we map coordinates to the [0,1000) range and add the special token notation <|box_start|><|box_end|> (Wang et al., 2024b). During Qwen2-VL-Base fine-tuning, we observe that Chinese datasets consistently degraded performance, possibly due to lower data quality or because evaluation benchmarks are primarily in English. Consequently, we use only English datasets for instruction tuning of Qwen2-VL-Base. InternVL2.5-Instruct, already possessing multilingual instruction-following capabilities, is fine-tuned using all available data.
Evaluation benchmark.
We carefully select specialized datasets to evaluate distinct abilities across tasks. (i) For VQA, we utilize VizWiz (Gurari et al., 2018) and GQA (Hudson & Manning, 2019) to assess general visual question answering proficiency. (ii) For Geometry, we incorporate multiple challenging subsets: “geometry reasoning”, “algebraic reasoning” and “geometry problem solving” from MathVista (Lu et al., 2024a), complemented by “metric geometry - angle”, “metric geometry - area”, “metric geometry - length” and “solid geometry” from MATH-Vision (Wang et al., 2024a). (iii) For Chart, we employ ChartQA (Masry et al., 2022), which tests reasoning and interpretation ability with charts and graphs. (iv) For OCR, our evaluation suite includes TextVQA (Singh et al., 2019) and OCRVQA (Mishra et al., 2019). (v) For Grounding, we implement referring expression comprehension using RefCOCO (Kazemzadeh et al., 2014), RefCOCO+ (Kazemzadeh et al., 2014), and RefCOCOg (Mao et al., 2016), which require models to identify specific objects in images based on natural language descriptions. All evaluation results are obtained using the VLMEvalKit (Duan et al., 2024) and LMMs-Eval (Zhang et al., 2024) libraries under the same settings to ensure fair comparison. When evaluating MathVista and MATH-Vision benchmarks, we utilize the GPT-4o-mini API to extract answers from the output.
For Omni-language models, we select audio-visual question answering task. This task requires multimodal understanding and spatio-temporal reasoning over audio-visual scenes. AVQA (Yang et al., 2022) targets real-world objects and activities. MUSIC-AVQA (Li et al., 2022) specifically focuses on musical performances.
Appendix D Discussions
D.1 Understanding the Task Vector
@@
WUDI Merging (Cheng et al., 2025) substitutes the transpose of the task vector $\bm{\tau}$ for the input $\bm{x}$. We reconsider the update process of the task vector, which can be formulated as follows:
$\bm{\tau}_{i,l}$ $=\textstyle\sum_{t=1}^{\mathrm{T}}-\eta\cdot\frac{\partial{\mathcal{L}(\bm{\theta}^{t-1}_{i})}}{\partial{\bm{\theta}^{t-1}_{i,l}}}$ (24)
$=\textstyle\sum_{t=1}^{\mathrm{T}}-\eta\sum^{N}_{n=1}\frac{\partial\mathcal{L}(\bm{\theta}^{t-1}_{i})}{\partial(\bm{\theta}^{t-1}_{i,l}\bm{x}_{i,l}^{n-1})}\cdot\frac{\partial(\bm{\theta}^{t-1}_{i,l}\bm{x}_{i,l}^{n-1})}{\partial\bm{\theta}^{t-1}_{i,l}}$ (25)
@@
$=\underbrace{\textstyle\sum_{t=1}^{\mathrm{T}}-\eta\sum^{N}_{n=1}\frac{\partial\mathcal{L}(\bm{\theta}^{t-1}_{i})}{\partial(\bm{\theta}^{t-1}_{i,l}\bm{x}_{i,l}^{n-1})}}_{\text{coefficient}}\cdot(\bm{x}_{i,l}^{n-1})^{\top},$ (26)
@@
where $\bm{\tau}_{i,l}$ denotes the task vector of task $i$ in linear layer $l$, and $\bm{\theta}^{t-1}_{i,l}$ represents the parameters of task $i$ in linear layer $l$ at time t1t-1. Each parameter in the linear layer can be interpreted as a weighted sum of input vectors across training iterations, with gradients serving as coefficients.
D.2 Limitation and Future Work
Due to resource constraints, our experiments were limited to models of 7B parameters. The public datasets we collected may contain lower-quality data. Future work will explore multilingual or reasoning-focused MLLM merging, incorporating visual chain-of-thought datasets (Yang et al., 2025; Li et al., 2025c) to support expert reasoning models. For evaluation, we plan to develop new benchmarks specifically designed to assess the reasoning capabilities of MLLMs.
D.3 Broader Impacts